2.3.53 Let g(x) = - x2 + 4x +1. Find and simplify g(-2). g(-2) = (Simplify your answer.)
Answer #1

2.3.53

\\\\g(x) = -x^2+4x +1 \\\\We \ have \ to \ determine \ g(-2) \\\\Putting \ in \ place \ of \ x \ we \ have \\\\g(-2) =-(-2)^2 +4*(-2)+1 \\\\g(-2) = -4 -8+1 \\\\g(-2) = -12+1 \\\\g(-2) = -11

The answer is g(-2) = -11